3. High-Low Method

Ziegler Inc. has decided to use the high-low method to estimate the total cost and the fixed and variable cost components of the total cost. The data for various levels of production are as follows:

Units Produced Total Costs
7,260 $566,720
2,860 364,320
4,660 540,000

a. Determine the variable cost per unit and the total fixed cost.

Variable cost: (Round to the nearest dollar.) $ ____ per unit
Total fixed cost: $______

b. Based on part (a), estimate the total cost for 3,640 units of production.

Total cost for 3,640 units: $______
Answer #1

A) Variable cost = (Cost at highest activity - Cost at lowest activity)/(Highest activity-Lowest activity)

= (566,720-364,320)/(7260-2860)

Variable cost = 46

Total fixed cost = 566,720 - (7260*46)

Total fixed cost = 232,760

B) Total cost for 3640 units

= (3640*46) + 232,760

= $400,200